------------------------------------------------------------
Date: Sat Oct 27 00:10:55 2001, Shift: 16:00-24:00
Supervisor: DO, with: 
Log book#10 pages 59 - 60
Shift Summary:
------------------------------------------------------------
1.Machine status:
  After a couple of unsuccessful ramps, MCR could store
  the beam from 19:00 to 20:30. They dumped for a RHIC 
  cryo ring access. They completed the access at 21:15    
  and the next store (beam cogged) started around 22:10. 
  Still running.
2.Achievements:
            Trig 1     Trig 6
  Run 5525 stopped after 10mn: trigger 7 downs scale was 1
  Run 5526    261k        84k  
  Run 5527    267k        85k
  Run 5528 still running (77k - 25k)  
3.Detector problems:
  The only bad thing I could notice is the absence
  of signals in TDC TOFW Top tube 92. Otherwise, things
  are in a good shape.  
------------------------------------------------------------
This archive was generated by hypermail 2b30 : Sat Oct 27 2001 - 00:12:13 EDT